# TempoFormer: A Transformer for Temporally-aware Representations in Change Detection

### v1

@@ 4.2 Baselines and Experimental Setup
We select baselines that are both post-level and stream-level (current post and recent history window, see §3.1).
Post-level: Random: post classification based on class distribution. BERT/RoBERTa: BERT (Devlin et al., 2018) or RoBERTa (Liu et al., 2019) fine-tuned using focal loss (Lin et al., 2017).
Llama2-7B-U (5/10-shot): Llama2-7B-chat-hf LLM (Touvron et al., 2023) with a crafted prompt for each dataset and few-shot (5/10) examples based on the distribution of input text, randomly sampled from each label distribution of each dataset, following Min et al. (2022). MistralInst2-7B-U(5/10-shot): MistralInst2-7B-U (Jiang et al., 2023): LLM with few shot examples. Same setting as for Llama2-7B-chat.
Stream-level: FFN History: Feed-forward on the concatenation of SBERT (Reimers and Gurevych, 2019) embeddings of the current post and the average of the window stream. SWNU (Tseriotou et al., 2023): feeds convolution-informed expanding path signature streams of SBERT representations and time into a BiLSTM modeling information progression. Seq-Sig-Net (Tseriotou et al., 2023): Sequential Network of SWNU modeled through a BiLSTM for long-term dependencies. BiLSTM: with a single layer on window stream SBERT embeddings. Llama2-7B-S (5-shot): Few-shot Llama2-7B-chat-hf prompted with the recent history of window 5 provided with each shot for context. MistralInst2-7B-S (5/10-shot): Same context-based setting as for Llama2-7B-S using Mistral-7B-Instruct-v0.2.
Evaluation: In line with published literature we report F1 scores for model performance, per class and macro-averaged. For each dataset we perform 5-fold cross validation with train/dev/test sets consisting of different timelines. We run and report the performance of each model on the exact same four random seeds (0,1,12,123) and report the average result on the test set. Appendix D provides information about implementation details and hyperparameter search.
@@ 5.1 Comparison against baselines
We present results for TempoFormer and baselines in Table 2. TempoFormer is the most performant in all three tasks based on macro-averaged F1. We note that recurrent models based on pre-trained BERT representations (BiLSTM for LRS and Topic Shift MI and Seq-Sig-Net for TalkLife), ranked second best. These models have been the SOTA for these datasets (Tseriotou et al., 2023, 2024). While the datasets are of different sizes, temporal characteristics, timeline length and change event distribution (see Table 1), TempoFormer retains its high performance, showcasing its generalisability for real-time change detection. Importantly, our model has the highest F1 for all minority classes, with the exception of Topic Shift MI, where other baselines have higher class-specific F1 scores for M but much lower F1 for R. Since TempoFormer operates on a contextual window of recent posts we select the appropriate window for each stream based on a window analysis, reported in §5.2.
We distinguish baselines into post and stream-level ones, noticing that smaller fine-tuned Language Models, even as simple as an FFN, allowing for stream-level context, score consistently better than post-level ones - with the exception of RoBERTa for TalkLife. This consistent finding underscores the importance of developing contextually informed representations for change detection. Few-shot prompted LLMs have consistently lower performance than smaller fine-tuned LMs, in line with reported poor temporal abilities of LLMs (Jain et al., 2023; Bian et al., 2023). For post-level, while Mistral’s performance improved from 5 to 10-shot, it is still barely above the random baseline and significantly behind BERT and RoBERTa. For LRS and Topic Shift MI the stream-level 5 and 10-shot Mistral performance increases, but falls way short of BERT/RoBERTa and all the stream-level models, indicating that although sequential context is important it is not modelled appropriately using LLMs. In line with (Wenzel and Jatowt, 2024), Llama2 suffers from generating responses outside the predefined classes, resulting in very low performance. TempoFormer demonstrates a generalisable architecture that enhances word-level post representations given the context, while modelling effectively the interplay between linguistic and temporal dynamics.
LRS TalkLife Topic Shift MI Model N-Sw Sw macro-avg IE IS O macro-avg M R macro-avg Post-level Random 61.4 37.5 49.5 11.2 4.5 84.4 33.4 35.9 63.9 49.9 Llama2-7B-U (5-shot) 22.4 50.6 36.5 10.1 7.5 31.9 16.5 46.6 45.4 46.0 MistralInst2-7B-U (5-shot) 71.4 28.0 49.7 23.3 4.1 67.8 31.7 46.4 44.6 45.5 Llama2-7B-U (10-shot) 8.8 52.5 30.7 12.8 6.2 31.3 16.7 48.5 39.5 44.0 MistralInst2-7B-U (10-shot) 71.2 30.5 50.8 27.6 3.5 72.1 34.4 42.6 55.7 49.1 BERT 69.0 45.3 57.1 43.9 28.1 86.8 52.9 36.0 70.0 53.0 RoBERTa 68.2 46.4 57.3 46.3 30.4 86.6 54.4 34.5 70.2 52.4 Stream-level FFN History 71.6 52.8 62.2 45.4 27.1 88.0 53.5 39.4 70.1 54.8 SWNU 75.5 55.5 65.5 48.0 29.3 89.5 55.6 38.7 66.0 52.3 Seq-Sig-Net 74.7 58.9 66.8 48.4 30.2 89.5 56.0 37.4 66.7 52.1 BiLSTM 75.0 60.7 67.8 46.1 27.0 89.2 54.1 37.8 73.8 55.8 Llama2-7B-S (5-shot) 2.2 50.2 26.2 15.5 7.6 24.2 15.7 52.6 1.3 27.0 MistralInst2-7B-S (5-shot) 58.3 50.2 54.3 22.0 4.6 70.0 32.2 42.3 57.3 49.8 MistralInst2-7B-S (10-shot) 54.4 51.8 53.1 23.4 3.5 74.9 33.9 37.8 63.7 50.8 TempoFormer (ours) 75.9 62.0 68.9 50.0 32.4 88.8 57.1 41.6 70.7 56.1
@@ 5.3 Ablations Study
In Table 3 we present an ablation study to assess each of TempoFormer’s components.
Temporal Rotary Multi-head Attention (MHA): By using the vanilla sequential distance version of RoPE in Multi-head attention instead of the temporal one, for the timestamped datasets, we see a drop in performance. This showcases the advantage of modeling linguistic streams while accounting for their temporal dynamics and the success of temporally distant RoPE. The relatively small drop in performance is due to the secondary role of temporal dynamics compared to the linguistic evolution in change detection.
@@
Stream embeddings: Removing these embeddings from TempoFormer layers brings noticeable performance drops in all datasets, showcasing the importance of propagating position information in building stream-aware and context-enhanced post embeddings. Topic Shift MI shows the largest drop of -1.4% among all its ablated models. Since this dataset does not obtain sequential signal from temporal dynamics, it relies on stream embeddings to model the distance between consecutive posts.
@@
LRS TalkLife Topic Shift MI Models N-Sw Sw macro-avg IE IS O macro-avg M R macro-avg TempoFormer 75.9 62.0 68.9 50.0 32.4 88.8 57.1 41.6 70.7 56.1 $\neg$¬Temporal RoPE 75.5 62.0 68.7 49.3 31.7 88.7 56.6 - - - $\neg$¬RoPE MHA 74.1 57.9 66.0 48.0 31.5 88.2 55.9 39.6 71.4 55.5 $\neg$¬Stream embed. 75.4 59.0 67.2 49.4 31.7 89.2 56.8 38.9 70.5 54.7 $\neg$¬Gate&Norm 74.5 61.3 67.9 49.8 31.1 88.7 56.6 40.7 69.6 55.2
@@ 5.4 The curious case of recurrence
RoTempoFormer consistently outperforms RoBERT for all datasets. RoTempoFormer strikes the right balance between jointly modeling context-aware post representations and recurrence in stream dynamics. Only for LRS do recurrence-based models have a better performance than TempoFormer. To examine this phenomenon we measure the diversity of each dataset with respect to its content and report it in Table 5. We report the BERTScore (Zhang et al., 2019b) and Cosine similarity between SBERT pairs of representations as well as the Outlier metric (Larson et al., 2019; Stasaski et al., 2020) on SBERT which measures the Euclidean distance between the (unseen) posts in the test set and the mean training corpus across folds and seeds for all datasets. Thus we assess both the semantic diversity and test set diversity. Across all metrics we consistently see that TalkLife is the most and LRS the least diverse. We postulate that for more diverse datasets like TalkLife, RoBERT has a really low performance, while it performs much better on less diverse ones. This could be due to: 1) overfitting due to recurrence and 2) inability of RoBERT to jointly model diverse context-aware representations, while capturing their evolution. RoTempoFormer, maintains its high performance, striking a good balance between modeling the context-aware post-level and the timeline-level dynamics. Importantly, we show that TempoFormer can be used as the foundation for temporal representation learning in other architectures.
LRS TalkLife Topic Shift MI model N-Sw Sw macro-avg IE IS O macro-avg M R macro-avg TempoFormer 75.9 62.0 68.9 50.0 32.4 88.8 57.1 41.6 70.7 56.1 RoBERT 75.8 62.3 69.0 36.7 3.3 88.4 42.8 33.3 75.7 54.5 RoTempoFormer 76.2 63.6 69.9 47.1 27.5 88.3 54.3 36.6 73.2 54.9
@@ 6 Conclusion
We introduce TempoFormer, a transformer-based model for change detection operating on textual (timestamped) streams. Importantly we do so by avoiding recurrence, and only modifying the last two layers of the transformer. Furthermore, TempoFormer has the ability to model the temporal distance between textual units through a modification of rotary positional embeddings. The model achieves new SOTA, outperforming recurrent and LLM-based models on three different change detection tasks with datasets of varying temporal granularity and linguistic diversity, without loss in generalisability. We demonstrate its usability as a foundation model in other architectures, striking the right balance between word-level, post-level and stream-level linguistic and temporal dynamics. Lastly, we showcase its flexibility in terms of base model integration, further boosting stream-level performance on par with post-level gains.
@@
The performance of our model, TempoFormer, is demonstrated on three datasets: LRS, TalkLife and Topic Shift MI. The LRS dataset is based on the publicly available RumourEval 2017 dataset (Gorrell et al., 2019) for stance detection, while the Topic Shift MI dataset is also a publicly available dataset based on human to human open domain conversations. Since the TalkLife dataset contains sensitive and personal user data, the appropriate Ethics approval was received from the Institutional Review Board (IRB), followed by data anonymisation and appropriate sensitive data sharing procedures. Access to this dataset was granted and approved by TalkLife  [^3] https://www.talklife.com/ through licensing for research purposes associated with the corresponding submitted proposal. All examples in the paper are paraphrased. Models were built on a secure server with authorised user-only access, while the labeled TalkLife dataset and the developed models are not indented to be made publicly available in order avoid any potential risks of unintended use.
@@ Appendix D Experimental Details
Appendix E LLM Prompts
To construct Mistral classification prompts we follow the recommended classification prompts as per provided guidelines  [^4] https://docs.mistral.ai/guides/prompting_capabilities/#classification. For constructing the Llama prompts we experimented with multiple prompts per dataset and identified the ones with the most stable performance. For fairer performance assessment we apply post-processing in LLM predictions to bucket them in the corresponding classification class (e.g. if the LLM generates esc we mark it as an escalation). In Tables 7, 8, 9, 10 we provide our LLM prompts for the LRS dataset.

### v2

@@ 4.2 Baselines and Experimental Setup
We select classification baselines that are both post-level (current post only) and stream-level (recent window of chronologically ordered posts including the current post, see §3.1). To account for the class imbalance, we use focal loss (Lin et al., 2017) for all the fine-tuned models.
Post-level:
Random: post classification based on probabilities of class distributions.
BERT/RoBERTa: BERT (Devlin et al., 2019) or RoBERTa (Liu et al., 2019) fine-tuned.
Llama2-7B-U (5/10-shot): In-context learning with Llama2-7B-chat-hf LLM (Touvron et al., 2023) using a crafted prompt for each dataset. Experiments on both 5 and 10 few-shot examples were randomly sampled to reflect the distribution of the dataset, following Min et al. (2022).
MistralInst2-7B-U(5/10-shot): Same setting as for Llama2-7B-U, using the Mistral- 7B-Instruct-v0.2 (Jiang et al., 2023) LLM.
Stream-level:
FFN History: Feed-forward network of 2 hidden layers on the concatenation of SBERT (Reimers and Gurevych, 2019) embeddings a) of the current post and b) averaged over the window posts.
SWNU (Tseriotou et al., 2023): Expanding windows of path signatures applied over learnable dimensionally-reduced data streams of SBERT representations and time and fed into a BiLSTM to model the information progression.
Seq-Sig-Net (Tseriotou et al., 2023): Sequential BiLSTM Network of SWNU units that capture long-term dependencies concatenated with the current post’s SBERT representation.
BiLSTM: Bidirectional single-layer recurrent network applied on the stream of SBERT embeddings.
Llama2-7B-S (5-shot): 5-shot in-context learning following the same set up as in Llama2-7B-U but including the recent history of window 5 in each shot (for context) instead of only the current post.
MistralInst2-7B-S (5/10-shot): Same 5 and 10 few-shot setting as for Llama2-7B-S, using the Mistral-7B-Instruct-v0.2 LLM.
Evaluation: In line with published literature we report F1 scores for model performance, per class and macro-averaged. For each dataset we perform 5-fold cross validation with train/dev/test sets consisting of different timelines. We run and report the performance of each model on the exact same four random seeds (0,1,12,123) and report the average result (as well as the standard deviation on macro-average) on the test set. Appendix D provides information about implementation details and hyperparameter search.
@@ 5.1 Comparison against baselines
We present results for TempoFormer and baselines in Table 2. TempoFormer is the most performant in all three tasks based on macro-averaged F1. We note that recurrent models based on pre-trained BERT representations (BiLSTM for LRS and Topic Shift MI and Seq-Sig-Net for TalkLife), ranked second best. The latter models have been the SOTA for these datasets (Tseriotou et al., 2023, 2024). While the datasets are of different sizes, temporal characteristics, timeline length and change event distribution (see Table 1), TempoFormer retains its high performance, showcasing its generalisability for real-time change detection. Importantly, our model has the highest F1 for all minority classes, with the exception of Topic Shift MI, where other baselines have higher class-specific F1 scores for M but much lower F1 for R. Since TempoFormer operates on a contextual window of recent posts we select the appropriate window for each stream based on a window analysis, reported in §5.2.
LRS TalkLife Topic Shift MI Model N-Sw Sw macro-avg IE IS O macro-avg M R macro-avg Post-level Random 61.4 37.5 49.5±0.510 11.2 4.5 84.4 33.4±0.080 35.9 63.9 49.9±0.332 Llama2-7B-U (5-shot) 22.4 50.6 36.5±0.000 10.1 7.5 31.9 16.5±0.000 46.6 45.4 46.0±0.000 MistralInst2-7B-U (5-shot) 71.4 28.0 49.7±0.000 23.3 4.1 67.8 31.7±0.000 46.4 44.6 45.5±0.000 Llama2-7B-U (10-shot) 8.8 52.5 30.7±0.000 12.8 6.2 31.3 16.7±0.000 48.5 39.5 44.0±0.000 MistralInst2-7B-U (10-shot) 71.2 30.5 50.8±0.000 27.6 3.5 72.1 34.4±0.000 42.6 55.7 49.1±0.000 BERT 69.0 45.3 57.1±0.995 43.9 28.1 86.8 52.9±0.140 36.0 70.0 53.0±0.186 RoBERTa 68.2 46.4 57.3±1.280 46.3 30.4 86.6 54.4±0.321 34.5 70.2 52.4±0.266 Stream-level FFN History 71.6 52.8 62.2±0.915 45.4 27.1 88.0 53.5±0.372 39.4 70.1 54.8±0.448 SWNU 75.5 55.5 65.5±0.715 48.0 29.3 89.5 55.6±0.461 38.7 66.0 52.3±0.749 Seq-Sig-Net 74.7 58.9 66.8±0.487 48.4 30.2 89.5 56.0±0.219 37.4 66.7 52.1±0.977 BiLSTM 75.0 60.7 67.8±1.400 46.1 27.0 89.2 54.1±0.113 37.8 73.8 55.8±0.672 Llama2-7B-S (5-shot) 2.2 50.2 26.2±0.000 15.5 7.6 24.2 15.7±0.000 52.6 1.3 27.0±0.000 MistralInst2-7B-S (5-shot) 58.3 50.2 54.3±0.000 22.0 4.6 70.0 32.2±0.000 42.3 57.3 49.8±0.000 MistralInst2-7B-S (10-shot) 54.4 51.8 53.1±0.000 23.4 3.5 74.9 33.9±0.000 37.8 63.7 50.8±0.000 TempoFormer (ours) 75.9 62.0 68.9±1.409 50.0 32.4 88.8 57.1±0.352 41.6 70.7 56.1±0.463
@@
We distinguish baselines into post and stream-level ones, noticing that smaller fine-tuned Language Models, even as simple as an FFN, allowing for stream-level context, score consistently better than post-level ones - with the exception of RoBERTa for TalkLife. This consistent finding underscores the importance of developing contextually informed representations for change detection. Few-shot prompted LLMs have consistently lower performance than smaller fine-tuned LMs, in line with reported poor performance of LLMs on temporal tasks (Jain et al., 2023; Bian et al., 2024). For post-level, while Mistral’s performance improved from 5 to 10-shot, it is still barely above the random baseline and significantly behind BERT and RoBERTa. For LRS and Topic Shift MI the stream-level 5 and 10-shot Mistral performance increases, but falls way short of BERT/RoBERTa and all the stream-level models, indicating that although sequential context is important it is not modelled appropriately with current LLMs. In line with (Wenzel and Jatowt, 2024), Llama2 suffers from generating responses outside the predefined classes, resulting in very low performance. TempoFormer demonstrates a generalisable architecture that enhances word-level post representations given the context, while modelling effectively the interplay between linguistic and temporal dynamics.
@@ 5.3 Ablations Study
In Table 3 we present an ablation study to assess the effect of each of TempoFormer’s components.
Temporal Rotary Multi-head Attention (MHA): By using the vanilla sequential distance version of RoPE in Multi-head attention instead of the temporal one, for the timestamped datasets, we see a drop in performance. This showcases the advantage of modelling linguistic streams while accounting for their temporal dynamics and the success of temporally distant RoPE. The relatively small drop in performance is due to the secondary role of temporal dynamics compared to linguistic evolution in change detection tasks.
@@
Stream embeddings: Removing only the $s11$ embedding from the top layer results in performance drop, signifying the importance of re-integrating the absolute post position for context enhancement of word representations. Further ablating both of $s10$ and $s11$ embeddings from TempoFormer layers brings even more noticeable performance drops in all datasets, showcasing the overall significance of propagating sequence position information in building stream-aware and context-enhanced post embeddings. Topic Shift MI shows the largest drop of -1.4% among all its ablated models. Since this dataset does not obtain sequential signal from temporal dynamics, it relies on stream embeddings to model the distance between consecutive posts.
@@
LRS TalkLife Topic Shift MI Models N-Sw Sw macro-avg IE IS O macro-avg M R macro-avg TempoFormer 75.9 62.0 68.9 50.0 32.4 88.8 57.1 41.6 70.7 56.1 $\neg$¬Temporal RoPE 75.5 62.0 68.7 49.3 31.7 88.7 56.6 - - - $\neg$¬RoPE MHA 74.1 57.9 66.0 48.0 31.5 88.2 55.9 39.6 71.4 55.5 $\neg$¬Stream embed. $s11$ 75.7 60.1 67.9 49.7 32.1 88.9 56.9 43.7 68.2 55.9 $\neg$¬Stream embed. $s10,s11$ 75.4 59.0 67.2 49.4 31.7 89.2 56.8 38.9 70.5 54.7 $\neg$¬Gate&Norm 74.5 61.3 67.9 49.8 31.1 88.7 56.6 40.7 69.6 55.2
@@ 5.4 The curious case of recurrence
LRS TalkLife Topic Shift MI model N-Sw Sw macro-avg IE IS O macro-avg M R macro-avg TempoFormer 75.9 62.0 68.9±1.409 50.0 32.4 88.8 57.1±0.352 41.6 70.7 56.1±0.463 RoBERT 75.8 62.3 69.0±0.689 36.7 3.3 88.4 42.8±0.565 33.3 75.7 54.5±0.303 RoTempoFormer 76.2 63.6 69.9±0.397 47.1 27.5 88.3 54.3±0.266 36.6 73.2 54.9±0.234
@@
Parameters Mean Train Time/Fold (min) model (million) LRS TalkLife Topic Shift RoBERT 110 14.9 36.0 97.8 TempoFormer 144 15.6 38.0 99.1 RoTempoFormer 145 15.5 37.4 98.9
Table 6: Model size and training time requirements for recurrent Transformer-based Models. Time experiments are averaged across all folds, epochs and seeds.
RoTempoFormer consistently outperforms RoBERT for all datasets. RoTempoFormer strikes the right balance between jointly modelling context-aware post representations and recurrence in stream dynamics. Only for LRS do recurrence-based models have a better performance than TempoFormer. To examine this phenomenon we measure the diversity of each dataset with respect to its content and report it in Table 5. We report the BERTScore (Zhang et al., 2019b) and Cosine similarity between SBERT pairs of representations as well as the Outlier metric (Larson et al., 2019; Stasaski et al., 2020) on SBERT which measures the Euclidean distance between the (unseen) posts in the test set and the mean training corpus across folds and seeds for all datasets. Thus we assess both the semantic diversity and test set diversity. Across all metrics we consistently see that TalkLife is the most and LRS the least diverse. We postulate that for more diverse datasets like TalkLife, RoBERT has a really low performance, while it performs much better on less diverse ones. This could be due to: 1) overfitting due to recurrence and 2) inability of RoBERT to jointly model diverse context-aware representations, while capturing their evolution. RoTempoFormer, maintains its high performance, striking a good balance between modelling the context-aware post-level and the timeline-level dynamics. Importantly, we thus show that TempoFormer can be used as the foundation for temporal representation learning in other architectures.
We further present the parameter and time requirements for the recurrent Transformer-based architectures (of Table 4) in Table 6. While both TempoFormer and RoTempoFormer require around 30% more parameters than RoBERT for training, this increase in model size is not prohibitive given the performance gains. While there is an increased model size, the overall computation requirements of less than 150M parameters are still low. Additionally, the mean training time for the TempoFormer family of models is only 1-6% higher than for RoBERT. Time requirements across all modelas are mainly dependent on the utterance length and chosen window size for each of the datasets.
@@ 6 Conclusion
We introduce TempoFormer, a transformer-based model for change detection operating on textual (timestamped) streams. Importantly we do so by avoiding recurrence, and only modifying the last two layers of the transformer. Furthermore, TempoFormer has the ability to model the temporal distance between textual units through a modification of rotary positional embeddings. The model achieves new SOTA, outperforming recurrent and LLM-based models on three different change detection tasks with datasets of varying temporal granularity and linguistic diversity, without loss in generalisability. We demonstrate its usability as a foundation model in other architectures, showing it strikes the right balance between word-level, post-level and stream-level linguistic and temporal dynamics. Lastly, we showcase its flexibility in terms of base model integration, further boosting stream-level performance on par with post-level gains.
@@
The performance of our model, TempoFormer, is demonstrated on three datasets: LRS, TalkLife and Topic Shift MI. The LRS dataset is based on the publicly available RumourEval 2017 dataset (Gorrell et al., 2019) for stance detection, while the Topic Shift MI dataset is also a publicly available dataset based on human to human open domain conversations. Since the TalkLife dataset contains sensitive and personal user data, the appropriate Ethics approval was received from the Institutional Review Board (IRB), followed by data anonymisation and appropriate sensitive data sharing procedures. Access to this dataset was granted and approved by TalkLife  [^3] https://www.talklife.com/ through licensing for research purposes associated with the corresponding submitted proposal. All examples in the paper are paraphrased. Models were built on a secure server with authorised user-only access. The labeled TalkLife dataset and the developed models are not intended for public release in order avoid potential risks of unintended use.
Acknowledgements
This work was supported by a UKRI/EPSRC Turing AI Fellowship (grant no. EP/V030302/1) and Keystone grant funding from Responsible Ai UK to Maria Liakata (grant no. EP/Y009800/1), the Alan Turing Institute (grant no. EP/N510129/1), and a DeepMind PhD Scholarship to Talia Tseriotou. The work was carried out while Adam Tsakalidis was employed by Queen Mary University of London. The authors would like to thank Jenny Chim, Dimitris Gkoumas and the anonymous reviewers for their valuable feedback.
@@ Appendix D Experimental Details
Appendix E Window Results
Full results for the window analysis are presented in Table 8.
Appendix F Dataset Examples
Here we provide a linguistic stream example from each dataset in Tables 9, 10, 11.
Table 9: LRS 12-utterance long stream example with labels
LRS Stream
Stream History:
U1 Approximately 50 hostages may be held captive at #Lindt café – local reports http://t.co/1ZlzKDjvSf #sydneysiege http://t.co/NvLr5kyQG8
L1 No Switch (support)
U2 @RT_com That’s an exaggeration, get your facts right.
L2 No Switch (deny)
U3 @RT_com I thought it was only 1 from the beginning
L3 No Switch (comment)
U4 @RT_com 50 Hostages now
L4 Switch (support)
U5 @RT_com they’re gonna fuck that dude up
L5 Switch (comment)
U6 @RT_com I pray for the safety of all the hostages; and that they are released soon.
L6 Switch (comment)
U7 @RT_com - "Approximately 50 hostages", in the article linked the first few lines says the number is closer to 13.
L7 No Switch (deny)
U8 Good thing Australia has strict gun laws. "@RT_com: Approximately 50 hostages may be held captive at #Lindt café http://t.co/1RFsbJWl7h
L8 No Switch (comment)
U9 @Simbad_Reb why don’t you get off Twitter and protect the next pre-school that will get hit by your infinite number of crazed gunmen
L9 No Switch (comment)
U10 @RT_com nah it’s 5000 or maybe 500. Or Whatever sounds more alarming
L10 Switch (deny)
U11 @RT_com dear God!!!
L11 No Switch (support)
Current Utterance:
U12 @NijatK There is a mental health problem not a gun problem.
L12 No Switch (comment)
Table 10: TalkLife 12-utterance long stream example with labels (paraphrased)
TalkLife Stream
Stream History:
U1 Going to a Taylor Swift concert last week is a blessing. I feel so empowered.
L1 None
U2 Shake it off, shake it off
L2 None
U3 I am really craving for this feeling of getting on stage, singing my own music. It really scares me and excites me at the same time but I want to give it a chance.
L3 None
U4 let me be brave enough to explore the unknown.
L4 None
U5 he couldn’t take his eyes off, what should I be thinking?
L5 None
U6 if someone makes intense eye contact would does this mean?
L6 None
U7 I feel the attraction but I won’t do anything to hurt him. I already hurt his feelings before.
L7 None
U8 Everyone pretends like it’s not a big deal, but I can’t get over the fact that I rushed my friend in the emergency room the other day. I’m deeply scarred and distressed.
L8 Switch (IS)
U9 I have been through so much trauma lately and I need to say it out loud that I feel broken
L9 Switch (IS)
U10 My inspiration for singing is a burning flame, right when I thought I lost it. All these experiences helped me to rediscover music, so grateful for everything
L10 None
U11 I’m struggling to get enough air. What’s happening to me?
L11 Switch (IS)
Current Utterance:
U12 Because if you want, I’ll take you in my arms and keep you sheltered, From all that I’ve done wrong
L12 None
Table 11: Topic Shift MI 12-utterance long stream example with labels, denoting speakers as A and B
Topic Shift MI Stream
Stream History:
U1/B what, what do you do, now?
L1 Major
U2/A Well, we have saved our newspapers for years and years because the, uh, Boy Scouts our boys have been involved in have, uh, had a huge recycling bin, over at Resurrection Lutheran Church
L2 Major
U3/B Uh-huh.
L3 Major
U4/A and, uh, so we’ve done that for quite some time,
L4 Major
U5/A but since the price of paper has gone down
L5 Major
U6/A like it’s about a fifth of what it used to be
L6 Major
U7/B Oh, really?
L7 Major
U8/A so the Boy Scout troop quit doing it when the City took it over.
L8 Major
U9/B Okay.
L9 Major
U10/A So now we just put ours out for the City of Plano.
L10 Major
U11/A Do you live in Plano?
L11 Rest
Current Utterance:
U12/B Yes,
L12 Rest
Appendix G LLM Prompts
To construct Mistral classification prompts we follow the recommended classification prompts as per provided guidelines  [^4] https://docs.mistral.ai/guides/prompting_capabilities/#classification. For constructing the Llama prompts we experimented with multiple prompts per dataset and identified the ones with the most stable performance. For fairer performance assessment we apply post-processing in LLM predictions to bucket them in the corresponding classification class (e.g. if the LLM generates esc we mark it as an escalation). In Tables 12, 13, 14, 15 we provide our LLM prompts for the LRS dataset.